Birch Company normally produces and sells 44.000 u
Answer #1

Qn no 1a.

Situation Normal Situation Production Continues during Strike Plant Shut Down
Sales unit 44000 13000 0
unit Sale price $             25.00 $         25.00 $             25.00
Unit Variable Cost $             21.00 $         21.00 $             21.00
Contribution per unit $              4.00 $           4.00 $               4.00
Total Contribution $     176,000.00 $   52,000.00 $                  -  
Fixed Cost $     219,000.00 $ 219,000.00 $     155,260.00
Profit $     (43,000.00) $(167,000.00) $    (155,260.00)
Profit during 2 months $     (86,000.00) $(334,000.00) $    (310,520.00)
Start up Cost $       14,000.00
Actual Profit during 2 months $     (86,000.00) $(334,000.00) $    (324,520.00)

So Net Income decresed by 324520 - 86000 = $238,520 in two months

1b.

If plant is not shut down and sales only 13000 units of RG-6 per month, Profit will be reduced by only (334000 - 86000)

= $ 248000, so Britch company should shut down the plant.

Answer : Yes

2.

If Plant is shut down:

Cost Saving Fixed Cost =( 60000+34000*11% ) * 2 - 14000 =$ 113480 during 2 months

Contribution of $113480 has to be earned from sale of 113480/4 = 28370 units during 2 months or 14185 units per month.

So Level of sales = 28370 units in two months